OpenStudy (anonymous):

Whats the equation for (-5,7),(-2,5)

jimthompson5910 (jim_thompson5910):

if you meant \[\Large y = -\frac{2}{3}x + \frac{11}{3}\], then you are correct KiwiBaby

jimthompson5910 (jim_thompson5910):

another way to express the equation is 2x + 3y = 11

OpenStudy (anonymous):

Whoa, did he delete my answer? What a douche..

jimthompson5910 (jim_thompson5910):

who? what do you mean?

OpenStudy (anonymous):

andres9823 the guy who asked the question.. My first response to the question is gone on my end.. looks like he deleted it. kind of rude..... Not you though, sorry for any misunderstanding.

jimthompson5910 (jim_thompson5910):

i have no idea of what message you meant so you're ok